# Automated Proof Generation for Rust Code via Self-Evolution

### v1

@@ Abstract
Ensuring correctness is crucial for code generation. Formal verification offers a definitive assurance of correctness, but demands substantial human effort in proof construction and hence raises a pressing need for automation. The primary obstacle lies in the severe lack of data — there is much less proof than code for LLMs to train upon. In this paper, we introduce SAFE, a novel framework that overcomes the lack of human-written proof to enable automated proof generation of Rust code. SAFE establishes a self-evolving cycle where data synthesis and fine-tuning collaborate to enhance the model capability, leveraging the definitive power of a symbolic verifier in telling correct proof from incorrect ones. SAFE also re-purposes the large number of synthesized incorrect proofs to train the self-debugging capability of the fine-tuned models, empowering them to fix incorrect proofs based on the verifier’s feedback. SAFE demonstrates superior efficiency and precision compared to GPT-4o. Through tens of thousands of synthesized proofs and the self-debugging mechanism, we improve the capability of open-source models, initially unacquainted with formal verification, to automatically write proof for Rust code. This advancement leads to a significant improvement in performance, achieving a 70.50% accuracy rate in a benchmark crafted by human experts, a significant leap over GPT-4o’s performance of 24.46%.
@@ 1 Introduction
Large Language Models (LLMs) have recently exhibited impressive capabilities in code generation (Roziere et al., 2023; Guo et al., 2024; Lozhkov et al., 2024; Google, 2024). However, the correctness of generated code cannot be guaranteed. To tackle this issue, prior research (Chen et al., 2022; Zhang et al., 2024a; Huang et al., 2023) has explored assessing generated code by test cases, which sometimes are also generated by LLMs. Although helpful, testing cannot cover all possible program inputs; they can reveal the presence of bugs but cannot prove their absence (Dahl et al., 1972). In contrast, formal verification employs a symbolic prover to mathematically prove that a program satisfies the desired properties under all possible inputs, without running the program. Unfortunately, to use formal verification, one needs to first formally express the desired properties, often in a special proof-oriented language, and then craft formal proofs, which takes substantial formal verification expertise (Zhang et al., 2024b). Therefore, to advance trustable code generation, automated formal verification, particularly automated proof generation, is in a pressing need.
@@
Two main approaches of proof automation have been explored. The first (Misu et al., 2024; Sun et al., 2024a; Pei et al., 2023; Liu et al., 2023; Chakraborty et al., 2023; Kamath et al., 2023; Chakraborty et al., 2024; Yao et al., 2023; Yang et al., 2024a) relies on well crafted prompt or in-context learning on LLMs. To compensate for the unfamiliarity of LLMs towards formal proof, this approach often requires intensive prompt engineering that hard codes proof-writing tips, many LLM invocations, and/or support from static program analysis, resulting in a limited generalization ability. The second approach (Polu et al., 2022; First et al., 2023; Azerbayev et al., 2023; Yang et al., 2024b) fine-tunes open-source LLMs. This approach has shown to be effective for Lean (Avigad, 2017) (for mathematical proving) and F (Swamy & Hrit, 2016) (a special proof-oriented language), where tens of thousands of human-written Lean and F proofs exist. However, how to apply this data-oriented approach to many other verification tools or proof languages that have much less exiting proof remains as an open question.
@@
In this paper, we tackle this open question in automating proof generation for Verus (Lattuada et al., 2023), the state-of-the-art verification tool for code written in Rust. Verus is a perfect target for us, since (1) it is one of the rare verification tools that can directly prove the correctness of code written in a popular language (Rust), which is increasingly widely used in production as a safer and better alternative to C/C++ (ONCD, 2024; Rivera et al., 2021); and (2) due to its short history, there are fewer than 500 Rust files currently verified by Verus on GitHub, too few for fine-tuning.
@@
The challenge of lacking data for Verus proof generation shows up at several levels. First, lack of suitable programs to quickly generate proof for. Not all Rust features are supported by Verus yet. Furthermore, large Rust projects typically require code refactoring to get verified (Zhou et al., 2024; Sun et al., 2024b), and yet small Rust programs may not have meaningful properties to prove (e.g., a hello-world function). Second, lack of Verus specifications. Before proving any Rust function, Verus needs a formal specification that describes the expected behavior of the function, such as the pre-condition (Line 7–9) and post-condition (Line 10–13) shown in Listing 1. Not surprisingly, these specifications do not exist except for in those existing verified Rust files on Github. Third, lack of Verus proofs. It took many months to write a couple of thousand lines of Verus proofs even for experts (Zhou et al., 2024; Sun et al., 2024b). It took more than ten years’ effort to accumulate those fine-tuning training data for Lean (Yang et al., 2024b) and F-Star (Chakraborty et al., 2024).
@@
In this paper, we propose the Self-evolving Automated prooF gEneration (SAFE) framework that overcomes the challenge of multi-level data scarcity. First, to quickly obtain a large number of proof-friendly Rust programs, we leverage GPT-4o to “translate” tens of thousands of small Python and Rust programs in popular code-synthesis datasets to Rust programs written in Verus-supported Rust features. Any program that cannot be compiled by Verus compiler is dropped.
@@
Second, to obtain formal specifications for these tens of thousands of Rust programs, SAFE uses a procedure of self-evolving specification generation. This procedure is bootstrapped by GPT-4o that is prompted to generate Verus-style specifications based on each Rust function and its associated natural language doc-string. In every following round, those specifications generated in previous rounds are used to train a new fine-tuned open-source LLM that produces more specifications in generally higher quality for the next round of finetuning. Keys to the success of this step include (1) using a quantitative metric (Lahiri, 2024) to differentiate high-quality specifications from low-quality ones, and only use the former for finetuning; and (2) the observation that we only need reasonably well, instead of perfect, specifications to enable proof generation in the next step.
@@
Next, to obtain tens of thousands of verified Rust programs, SAFE similarly uses a self-evolving procedure. Proof is much harder to synthesize than specifications. Even with careful prompt engineering, GPT-4o only managed to synthesize correct proof for $<20\%$ of programs in the dataset with specifications synthesized above, after a month of non-stop invocation. Fortunately, this is sufficient to bootstrap SAFE. Through rounds of self-evolving, the quantity and quality of synthesized proof keeps increasing, while the fine-tuned open-source model’s capability of proof synthesis keeps getting augmented. These models also generate proof much faster than GPT-4o. The keys to the success here include (1) the ability of Verus in authoritatively and quickly telling correct proof from incorrect ones, which allows SAFE to sift through the huge amount of low-quality data in early rounds without any human labeling, and (2) the reasonable quality of SAFE-synthesized specifications that allow SAFE to largely avoid trivial proofs, which have little usage in fine-tuning.
@@
Finally, SAFE re-purposes the huge number of incorrect proofs generated in the previous step to add self-debugging capability into its models. In each round, whenever a correct proof $P_{\checkmark}$ is synthesized after several attempts, the triplet of an earlier incorrect proof $P_{X}$, the verification error reported by Verus on $P_{X}$, and $P_{\checkmark}$, becomes a training data point, which fine-tunes the model’s capability in debugging and repairing proof— a great side effect of having incapable models early on.
@@
In our experiments, SAFE leverages DeepSeekCoder (Guo et al., 2024) as the generator, successfully synthesizing 19,017 formal specifications and 9,706 verified Rust functions, from a dataset comprising 45,395 Rust functions sourced from the MBPP (Austin et al., 2021) training split and the CodeNet (Puri et al., 2021) training dataset. Note that, the initial Rust dataset contains zero lines of formal specification or proof. Our evaluation on a human-curated Verus benchmark with human-written specification, VerusBench, and a synthetic benchmark, CodeNet-Test demonstrate that SAFE empowers DeepSeekCoder, which is initially unacquainted with Verus, to achieve 43.17% and 43.83% accuracy on the two benchmarks by direct generation, far surpassing GPT-4o’s performance of 11.51% and 0.28%. Furthermore, the model’s accuracy reaches 79.14% in VerusBench and 48.43% in CodeNet-Test once its self-debugging feature is used.
We believe a key to SAFE’s self-evolving framework lies in its automated measuring and filtering mechanisms for every step of data synthesis — Verus compiler for the first step, the quantitative specification-quality metric for the second step, and Verus verification for the last two steps. These mechanisms allow high-quality training data to gradually get accumulated without human labelling. We believe this SAFE approach can be extended and adapted to enhance any LLMs and other formal languages that face data scarcity issues, offering a promising solution for proof automation. Finally, we will open source our code and benchmarks.
@@ 2 Background and Related Work
1use vstd::prelude::*;
2fn main() {}
3
4verus!{
6fn binary_search(v: &Vec<u64>, k: u64) -> (r: usize)
@@ 2.1 Verus verification tool
To formally prove whether a Rust function is guaranteed to satisfy its specification, Verus (Lattuada et al., 2023) statically analyzes the Rust code, the specification, and the proof annotations, and forms queries for the underlying SMT-solver (e.g., Z3 (De Moura & Bjørner, 2008)) to formally solve. By leveraging the type system in Rust and allowing both specification and proof annotations to write in Rust syntax, Verus has become the state-of-the-art verification tool for Rust, one of the most popular programming languages (JE & CT, 2020; Fulton et al., 2021), and has been used to successfully verify large Rust systems (Sun et al., 2024b; Zhou et al., 2024). Unfortunately, since Verus has been developed for only about three years, there are less than 500 Verus verified Rust files on Github so far. Meanwhile, although Verus specifications and proof are written mostly in Rust, they are still difficult to generate for both human beings (Sun et al., 2024b; Zhou et al., 2024) and LLMs.
Listing 1 shows a Rust function that implements binary search. At high level, the function specification (Line 7–13) states that the post-condition, expressed in a ensures code block, is guaranteed to hold as long as the pre-condition, in a requires block, is true. For this function, its pre-condition (Lines 7-9) states that the input array is ordered in an ascending manner and that the input value to be searched k must exist in the array (forall and exists are Verus quantifiers). The post-condition requires that the return value $r$ should be a valid index pointing to the input value k.
For simple programs/specifications, Verus can accomplish the formal verification without any extra annotations (we refer to these cases as trivial proof). Unfortunately, for functions and specifications that involve loops, collections, and quantifiers, Verus often needs users to provide proof annotations. The loop invariants specified on Line 18–21 are one type of proof annotations. They specify what properties are true right before and right after every loop iteration. Verus will prove the correctness of each loop invariant and then use proved loop invariants to help prove the function specification. If any loop invariant is incorrect or if needed invariants are missing, the proof will fail. For more complicated tasks, other types of proof annotations like assert and lemma functions may be needed.
@@ 2.2 Prior work in proof-benchmark building and self-evolving framework
There was much effort recently in building verification dataset of existing human written proofs (Loughridge et al., 2024; Zhang et al., 2024b; Chakraborty et al., 2024). Unfortunately, since writing formal proof takes expertise beyond normal coding, this dataset building approach is difficult to scale. Recent work has explored augmenting the Dafny-proof dataset using proofs synthesized by GPT-4 through few-shot learning and human-written proof examples (Misu et al., 2024; Sun et al., 2024a). However, due to the limited proof-generation power of GPT-4, the number of proofs in these dataset are fewer than 200, with many being trivial proofs.
Self-evolving style of learning has been explored before in other context. While (Gulcehre et al., 2023) propose a reinforced self-training framework in the machine translation task, (Polu et al., 2022) leverages the expert iteration strategy for mathematical theorem proving with Lean. Our work applies the idea of self-evolving to a very different task, synthesizing correctness proof for Rust code, with different challenges: unlike previous work, Verus does not have a large collection of existing data for bootstrapping; there is much less a concern of generating “specification” in mathematical theorem proving; furthermore, unlike interactive theorem provers that decompose a proof problem into small steps through tactics (e.g., Lean), Verus uses SMT-solver with hints from proof annotations to prove the whole function correctness as a whole and hence presents completely different challenges to proof synthesis.
@@ 3 Approach
As illustrated in Figure 1, SAFE involves two self-evolving (Tao et al., 2024; Gulcehre et al., 2023) procedures. The first procedure synthesizes specifications that are used as inputs to the second procedure, while the second procedure produces the end-result of SAFE — a fine-tuned LLM that can automatically synthesize proof for Rust code.
In Algorithm 1, we show the two procedures of our self-evolving framework. In both procedures, we use GPT-4o to generate the round-00 data $spec/proof\textit{-}data_{0}$. In each round $r$, we use data collected and filtered from rounds 0 to $r-1$ to fine-tune $model_{r}$ based on its preceding $model_{r-1}$. At the end of each round, $model_{r}$ is used to generate the round-$r$ data, which is then graded and filtered. High-quality data is preserved for the next round. Low-quality specifications are discarded; incorrect proofs are re-purposed to become self-debugging training data for the next round. This iterative approach substantially improves the volume and quality of the synthesized data.
In the following, we present more details for each step of SAFE.

### v2

@@ Abstract
Ensuring correctness is crucial for code generation. Formal verification offers a definitive assurance of correctness, but demands substantial human effort in proof construction and hence raises a pressing need for automation. The primary obstacle lies in the severe lack of data—there is much fewer proofs than code snippets for Large Language Models (LLMs) to train upon. In this paper, we introduce SAFE, a framework that overcomes the lack of human-written proofs to enable automated proof generation of Rust code. SAFE establishes a self-evolving cycle where data synthesis and fine-tuning collaborate to enhance the model capability, leveraging the definitive power of a symbolic verifier in telling correct proofs from incorrect ones. SAFE also re-purposes the large number of synthesized incorrect proofs to train the self-debugging capability of the fine-tuned models, empowering them to fix incorrect proofs based on the verifier’s feedback. SAFE demonstrates superior efficiency and precision compared to GPT-4o. Through tens of thousands of synthesized proofs and the self-debugging mechanism, we improve the capability of open-source models, initially unacquainted with formal verification, to automatically write proofs for Rust code. This advancement leads to a significant improvement in performance, achieving a 52.52% accuracy rate in a benchmark crafted by human experts, a significant leap over GPT-4o’s performance of 14.39%.
@@ 1 Introduction
Large Language Models (LLMs) have recently exhibited impressive capabilities in code generation (Roziere et al., 2023; Guo et al., 2024; Lozhkov et al., 2024; Google, 2024). However, the correctness of generated code cannot be guaranteed. To prove that a program satisfies the desired properties under all possible inputs, without running the program, we need formal verification. Unfortunately, formal verification is difficult to conduct, as one needs to first formally express the desired properties, often in a special proof-oriented language, and then craft formal proofs, taking substantial formal verification expertise (Zhang et al., 2024). Therefore, to advance trustable code generation, automated formal verification, particularly automated proof generation, is in a pressing need.
@@
Two main approaches of proof automation have been explored. The first (Misu et al., 2024; Sun et al., 2024a; Pei et al., 2023; Liu et al., 2023; Chakraborty et al., 2023; Kamath et al., 2023; Chakraborty et al., 2024; Yao et al., 2023; Yang et al., 2024a) relies on well crafted prompt or in-context learning on LLMs. To teach LLMs about formal proofs, this approach often requires intensive prompt engineering that hardcodes proof-writing tips, and/or requires support from static program analysis, resulting in a limited generalization ability. The second approach (Polu et al., 2022; First et al., 2023; Azerbayev et al., 2023; Yang et al., 2024b) fine-tunes open-source LLMs. This approach has been shown to be effective for Lean (Avigad, 2017) (for mathematical proving) and F (Swamy et al., 2016) (a special proof-oriented language), where tens of thousands of human-written Lean and F proofs exist. However, how to apply this data-oriented approach to many other verification tools or proof languages that have much fewer exiting proofs remains as an open question.
@@
In this paper, we tackle this open question for Verus (Lattuada et al., 2023), the state-of-the-art verification tool for code written in Rust. Verus is a perfect target for us, since (1) it is one of the rare verification tools that can directly prove the correctness of code written in a popular language (Rust), which is increasingly popular in production as a safer and better alternative to C/C++ (ONCD, 2024; Rivera et al., 2021); and (2) due to its short history, there are fewer than 500 Rust files currently verified by Verus on GitHub, too few for fine-tuning.
@@
The challenge of lacking data for Verus proof generation shows up at three main levels. First, lack of suitable programs to quickly generate proofs for. Not all Rust features are supported by Verus yet. Furthermore, large Rust projects typically require code refactoring to get verified (Zhou et al., 2024; Sun et al., 2024b), and yet small Rust programs may not have meaningful properties to prove (e.g., a hello-world function). Second, lack of Verus specifications. Before proving any Rust function, Verus needs a formal specification that describes the expected behavior of the function, such as the pre-condition (Lines 7–9) and post-condition (Lines 10–13) shown in Listing 1. Not surprisingly, these specifications do not exist except for in those existing verified Rust files on GitHub. Third, lack of Verus proofs. It took many months to write a couple of thousand lines of Verus proofs even for experts (Zhou et al., 2024; Sun et al., 2024b). It took more than ten years’ effort to accumulate those fine-tuning training data for Lean (Yang et al., 2024b) and F-Star (Chakraborty et al., 2024).
@@
In this paper, we propose the Self-evolving Automated prooF gEneration (SAFE) framework that overcomes the challenge of multi-level data scarcity. First, to quickly obtain a large number of proof-friendly Rust programs, we leverage GPT-4o to “translate” tens of thousands of small Python and Rust programs in popular code-synthesis datasets to programs written in Verus-supported Rust syntax. Any program that cannot be compiled by the Verus compiler is dropped.
@@
Second, to obtain formal specifications for these tens of thousands of Rust programs, SAFE uses a few rounds of self-evolving specification generation. The bootstrapping round prompts GPT-4o to generate Verus-style specifications based on each Rust function and its associated natural language doc-string. In every following round, specifications generated in previous iterations are used to train a new fine-tuned open-source LLM that produces more specifications in generally higher quality for the next round of finetuning. Keys to the success of this step include (1) using a quantitative metric (Lahiri, 2024) to differentiate high-quality specifications from low-quality ones, and use the former for only finetuning; and (2) the observation that we need only reasonably well, instead of perfect, specifications to enable proof generation in the next step.
@@
Next, to obtain tens of thousands of verified Rust programs, SAFE similarly uses a self-evolving procedure. Code proofs are very difficult to synthesize: even with careful prompt engineering, GPT-4o managed to synthesize correct proofs for only $<20\%$ of programs in the the preceding dataset with synthesized specifications, after a month of non-stop invocations. Fortunately, this is sufficient to bootstrap SAFE. Through rounds of self-evolving, the quantity and quality of synthesized proofs keep increasing, while fine-tuned open-source models’ capability of proof synthesis keeps getting augmented. These models also generate proofs much faster than GPT-4o. The keys to the success here include (1) the ability of Verus in authoritatively and quickly telling correct proofs from incorrect ones, allowing SAFE to sift through the huge amount of low-quality data in early rounds without any human labeling, and (2) the reasonable quality of SAFE-synthesized specifications that allow SAFE to largely avoid trivial proofs, which have little usage in fine-tuning.
@@
Finally, SAFE re-purposes the huge number of incorrect proofs generated in the previous step to add the self-debugging capability into its model. In each round, when a correct proof $P_{\checkmark}$ is synthesized after several attempts, the triplet of an earlier incorrect proof $P_{X}$, the verification error reported by Verus on $P_{X}$, and $P_{\checkmark}$, becomes a training data point, which fine-tunes the model’s capability in debugging and repairing proof—a great side effect of having incapable models early on.
@@
In our experiments, SAFE leverages DeepSeekCoder (Guo et al., 2024) as the generator, successfully synthesizing 19,017 formal specifications and 9,706 verified Rust functions, from a dataset comprising 45,395 Rust functions sourced from the MBPP (Austin et al., 2021) training split and the CodeNet (Puri et al., 2021) training dataset. Note that the initial Rust dataset contains zero lines of formal specification or proof. Our evaluation on a human-curated Verus benchmark with human-written specifications, VerusBench, and a synthetic benchmark, CodeNet-Test, demonstrates that SAFE empowers DeepSeekCoder, which is initially unacquainted with Verus, to achieve 43.17% and 43.83% accuracy on the two benchmarks by direct generation, far surpassing GPT-4o’s performance of 11.51% and 0.28%, respectively. Furthermore, the model’s accuracy reaches 79.14% in VerusBench and 48.43% in CodeNet-Test once its self-debugging feature is used.
@@ 2 Background and Related Work
1verus!{
2// Performs a binary search on a sorted vector of 64-bit unsigned integers (u64) to find the index of a given target value.
3fn binary_search(v: &Vec<u64>, k: u64) -> (r: usize)
@@ 2.1 Verus verification tool
Verus (Lattuada et al., 2023) statically analyzes every proof target (i.e., a Rust function and its specification) and any given proof annotations, and forms queries for the underlying SMT solver (e.g., Z3 (De Moura & Bjørner, 2008)) to solve. By leveraging the type system in Rust and allowing both specification and proof annotations to be written in Rust syntax, Verus has become the state-of-the-art verification tool for Rust, one of the most popular programming languages (Perkel, 2020; Fulton et al., 2021), and has been used to successfully verify large systems (Sun et al., 2024b; Zhou et al., 2024). Unfortunately, since Verus has been developed for only about three years, there are fewer than 500 Verus verified Rust files on GitHub.
Listing 1 shows a Rust function that implements binary search. Its specification includes a pre-condition enclosed in a requires block and a post-condition enclosed in an ensures block (Lines 4-10). The pre-condition states that the input array is ordered in an ascending way and the input value k to be searched exists in the array (forall and exists are Verus quantifiers); the post-condition requires that the return value $r$ should be a valid index pointing to the input value k.
For simple programs/specifications, Verus can accomplish the formal verification without any extra annotations (we refer to these cases as trivial proofs). Unfortunately, for functions and specifications that involve loops, collections, and quantifiers, Verus often needs users to provide proof annotations. The loop invariants specified on Lines 16–18 are one type of proof annotations. They specify what properties are true right before and right after every loop iteration. Verus will prove the correctness of each loop invariant and then use proved loop invariants to help prove the function specification. If any loop invariant is incorrect or if needed invariants are missing, the proof will fail. For more complicated tasks, other types of proof annotations like assert and lemma functions may be needed.
@@ 2.2 Prior work in proof-benchmark building and self-evolving framework
There has been much effort recently in building datasets of existing human written proofs (Loughridge et al., 2024; Zhang et al., 2024; Chakraborty et al., 2024). Unfortunately, since writing proofs takes expertise beyond normal coding, this dataset-building approach is difficult to scale. Recent work has explored augmenting the Dafny-proof dataset using proofs synthesized by GPT-4 through few-shot learning and human-written proof examples (Misu et al., 2024; Sun et al., 2024a). However, due to the limited proof-generation power of GPT-4, the number of proofs in these datasets are fewer than 200, with many being trivial proofs.
Self-evolving style of learning has been explored in other contexts: A reinforced self-training framework is proposed in the machine translation task (Gulcehre et al., 2023), and the expert iteration strategy is leveraged for math proving with Lean (Polu et al., 2022). Our work applies self-evolving and expert iteration to synthesize proofs for Rust code. This new task raises different challenges from prior tasks like writing math proofs (Polu et al., 2022) and hence requires different designs.
One obvious challenge is the issue of data scarcity discussed in Section 1: SAFE does not have access to billions of tokens of manually-written examples (e.g., math proofs) and hence cannot rely on fine-tuned models for bootstrapping as in prior work; in fact, SAFE does not even have access to a large quantity of proof problems and hence has to synthesize problems by itself (i.e., Verus-compatible Rust functions and the associated specifications).
Another major challenge is related to the underlying verification engine. Supported by an interactive theorem prover, Lean (Avigad, 2017), math-proof synthesis is naturally decomposed to many small steps or tactics, with clear judgment about what are useful intermediate proofs. In contrast, like many code-verification tools (Leino, 2010; Swamy et al., 2016), Verus leverages a SMT solver to prove the correctness of a function as a whole, while proof annotations are used as hints to the SMT solver. It is difficult to decompose Verus-proof synthesis into small steps, and it is difficult to judge whether incorrect proof annotations are useful. Consequently, step-wise search strategies used in prior work do not apply here. Instead, whole-proof debugging is crucial for SAFE.
We believe that SAFE presents a new application of the self-evolving and expert iteration philosophy to an important and challenging task—code-proof automation. Through its automated measuring and filtering at every step of data synthesis—Verus compiler for program transpilation, the quantitative specification-quality metric for specification synthesis, and Verus verification for proof-synthesis/debugging, SAFE overcomes those unique challenges and allows high-quality training data to gradually accumulate without manual effort.
@@ 3 Approach
As illustrated in Figure 1, SAFE involves two self-evolving (Tao et al., 2024; Gulcehre et al., 2023) procedures. The first procedure synthesizes specifications that are used as inputs to the second procedure, while the second procedure produces the end-result of SAFE—a fine-tuned LLM that can automatically synthesize proofs for Rust code. As listed in Algorithm 1, both procedures use GPT-4o to generate the round-00 data $spec/proof\textit{-}data_{0}$. In each round $r$, we use data collected and filtered from earlier rounds, $10...r-$ 1, to fine-tune $\text{model}_{r}$ based on its preceding $\text{model}_{r-1}$. At the end of each round, $\text{model}_{r}$ is used to generate the round-$r$ data. High-quality specifications are preserved for the next round. Low-quality specifications are discarded; incorrect proofs are re-purposed to become self-debugging training data for the next round.